A load of 20 N is applied to the ends of a wire 4 m long, and produces an extension of 0.24 mm. If the diameter of the wire is 2 mm, calculate the stress on the wire, its strain, and the value of Young’s ,modulus.
1
Expert's answer
2020-10-27T08:23:18-0400

The stress on the wire "\sigma=\frac{F}{S} =\frac{20}{\pi\times\frac{(2\times 10^{-3})^2}{4}}=6.37 \times 10^6" PA.

Strain "\epsilon=\frac {\Delta L}{L}=\frac{0.24 \times 10^{-3}}{4}=6\times 10^{-5}" .

Young's modulus "E=\frac{\sigma}{\epsilon}=\frac{6.37\times 10^6}{6\times 10^{-5}}=1.06\times 10^{11}" PA.